9. In the expression (1+x5+x9)10 find the coefficient of a) x23 b) x32

1
Expert's answer
2022-02-24T15:40:44-0500

In multinomial expansion of "(1+x^5+x^9)^{10}," every term will be of form:


"C(10; m, n, p)\times1^m\times(x^5)^n\times (x^9)^p,"

where "m + n + p =10."

a) For the coefficient of "x^{23}," "(5n + 9p)" should be equal to "23." There is only one pair exist for this condition to hold i.e "(1,2)."

"(m,n,p)" will be "(7, 1, 2)."


"C(10; 7, 1, 2)=\dfrac{10!}{7!1!2!}=\dfrac{10(9)(8)}{2}=360"

The coefficient of "x^{23}" is "360."


b) For the coefficient of "x^{32}," "(5n + 9p)" should be equal to "32." There is only one pair exist for this condition to hold i.e "(1,3)."

"(m,n,p)" will be "(6, 1, 3)."

"C(10; 6, 1, 3)=\dfrac{10!}{6!1!3!}=\dfrac{10(9)(8)(7)}{6}=840"

The coefficient of "x^{32}" is "840."